# Stability of Non-Hyperbolic Sets

### v1

@@ 6.2 Newhouse phenomenon: infinitely many sinks
For the last property, let $(t^{*},0)$ such that ${\log{\lambda(t^{*},0)^{-1}}}/{\log\mu(t^{*},0)}$ is the maximum of ${\log{\lambda(t,0)^{-1}}}/{\log\mu(t,0)}$. Consider a sequence of first generation Newhouse boxes $\mathcal{P}_{n,n_{0}}\in\mathfrak{N_{1}}$ accumulating at $(t^{*},0)$. This is possible because of Lemma 20. Choose $\epsilon>0$ and let $n$ be maximal such that $\epsilon\leq{\epsilon_{0}}/{\mu(t^{*},a^{*})^{2n}}$. Because (6.2), (6.3) and the fact that the vertical size of $\mathcal{P}_{n,n_{0}}$ is ${\epsilon_{0}}/{\mu(t^{*},a^{*})^{2n}}$, we need at least ${K}/{\lambda(t^{*},a^{*})^{\theta n}}$ balls of radius $\epsilon$ to cover $\overline{NH}\cap\mathcal{P}_{n,n_{0}}$. As consequence
@@
and $MD(NH)\geq{2}/{3}$, where we used (2.17). ∎
Remark 12.
@@ 7 Dynamical Partitions, Directed Graphs and Strips
In this section, we introduce the objects and the induction step (see Proposition 7) that will allow to build later the main player of this work: the lamination and the invariant Cantor set.
@@ 7.1 Dynamical Partitions
We give now the definition of a dynamical partition associated to a map. The reader can refer to the Figure 9. The definition models the combinatorics as described in Subsection 4.1.
@@
Recall the rescaling done in Subsection 2.1.1 while introducing our families of unfoldings and let $Q$ be the domain of linearization.
We denote by $W^{u}_{\text{loc}}(p)$ a connected component of $W^{u}(p)\cap Q$ and by $W^{s}_{\text{loc}}(p)$ a connected component of $W^{s}(p)\cap Q$.
Definition 7.
@@
    • $f\left(C^{N-1}\right)\subset Q\setminus f(Q)$,
    • $C^{0}\cap W^{u}_{\text{loc}}(p)\neq\emptyset$, where $W^{u}_{\text{loc}}(p)$ is a connected component of $W^{u}(p)\cap Q$,
    • $f\left(C^{N-1}\right)\cap W^{s}_{\text{loc}}(p)\neq\emptyset$.
@@
    • $f\left(H^{M-1}\right)\subset Q\setminus f(Q)$,
    • $H^{0}\cap W^{u}_{\text{loc}}(p)\neq\emptyset$, where $W^{u}_{\text{loc}}(p)$ is a connected component of $W^{u}(p)\cap Q$,
    • $f\left(H^{M-1}\right)\cap W^{s}_{\text{loc}}(p)\neq\emptyset$,
    • $f^{M}\left(H_{0}\cap W^{u}_{\text{loc}}(p)\right)$ intersects transversally $W^{u}_{\text{loc}}(p)$ in the point $q_{2}$.
Definition 8.
@@
where diam(U) is the diameter of $U$.
Moreover we denote by $\mathcal{D}$ the union of elements in $\mathfrak{D}$. Namely,
@@ 7.2 Directed Graphs
Remark 13.
Observe that $q$ has three ingoing and outgoing edges. All other vertices have exactly one ingoing and one outgoing edge.
Let $X$ be the $(1,N,M)$ directed graph. A loop $\lambda$ is a path in $X$ which starts and ends in $q$. The loops $\lambda^{q}$, $\lambda^{h}$ and $\lambda^{c}$ are called the prime loops of $X$, respectively $q$-loop, $h$-loop and $c$-loop.
Observe that a loop $\lambda$ is a concatenation of prime loops $\gamma_{i}$, namely
@@
Given $N,M,\tilde{N},\tilde{M}$ positive natural numbers, consider $the (1,N~,M~)(1,\tilde{N},\tilde{M}) directed graph X~\tilde{X}$ and the graph morphism
@@
In the following lemma we introduce the projection of $D$ on $X$.
Lemma 21.
@@ 7.3 Strips
We introduce strips in the plane that will enable us to later construct, inductively, the object of primary interest to us: the lamination. Observe that the strip are an abstract generalization of the $\mathcal{B}_{n}$ strips built earlier to contain the curves of secondary tangencies.
Definition 10.
@@
    $\left|s_{l}(t)-s_{u}(t)\right|\leq C\rho^{r(1+\alpha)-1},$
    where $0<\rho<1$, $\rho<\max_{t}\left\{\frac{1}{\mu(t)}\right\}$, $0<\alpha<1$ and $C$ is a positive constant.
  • $S$ contains $r$ global curves $\left\{b_{k}\right\}_{k=1}^{r}$, whcih are graph of functions $b_{k}:[-t_{0},t_{0}]\mapsto\mathbb{R}$ for all $k$, such that the following holds:
@@
We start by defining $r$-tangency strips, with $r$ a positive real number; see Figure 11.
@@ 7.4 The induction step
Proposition 7.
Let $S$ be a $r$-tangency strip and let $\left\{b,U\right\}$ be a $(1,N,M)$ descendant of $S$. Then there exist $n\in\mathbb{N}$, $n>>r$ large enough and a countable sequence of regions $\left\{S_{j}\right\}_{j=1}^{\infty}\subset U\subset S$ satisfying the following:
@@
    $\forall j$, $S_{j}$ is a $(n+j)$-tangency strip,
@@
Moreover, for all $j$ and for all $k\in\left\{1,\dots,n+j\right\}$ let $\left\{b_{k}^{(j)},U_{k}^{(j)}\right\}$ be a $\left(1,N_{k}^{(j)},M_{k}^{(j)}\right)$ descendant of $S_{j}$. Then the following holds:
@@
    $\exists C>0$ such that $\forall t\in\left[-t_{0},t_{0}\right]$,
    $\frac{1}{C}\frac{1}{\mu(t)^{n+j}}\leq\left|b_{k}^{(j)}(t)-b(t)\right|\leq\frac{C}{\mu(t)^{n+j}},$
@@
    $\exists\tilde{C}>0$ such that
    $\frac{1}{\tilde{C}}\frac{n+j}{\mu(t)^{n+1+j}}\frac{\partial\mu}{\partial t}\leq\left|\frac{db_{k}^{(j)}}{dt}-\frac{db}{dt}\right|\leq\tilde{C}\frac{n+j}{\mu(t)^{n+1+j}}\frac{\partial\mu}{\partial t},$ (7.1)
@@
    $\forall f\in U_{k}^{(j)}\subset U$, let $\mathfrak{D}^{(j)}_{k}(f)$ be the $\left(1,N_{k}^{j},M_{k}^{j}\right)$ dynamical partition of $f$ and let $\mathfrak{D}(f)$ be the $\left(1,N,M\right)$ dynamical partition of $f$. Then $\mathfrak{D}^{(j)}_{k}(f)$ is a refinement of $\mathfrak{D}(f)$ and
    $\text{mesh}\left(\mathfrak{D}^{(j)}_{k}(f)\right)<\left(\frac{1}{k}\right)\text{mesh}\left(\mathfrak{D}(f)\right).$
Furthermore, if $X$ is the $\left(1,N,M\right)$ directed graph, $X_{k}^{j}$ is the $\left(1,N_{k}^{j},M_{k}^{j}\right)$ directed graph, $i:\mathcal{D}_{k}^{(j)}\mapsto\mathcal{D}$ is the inclusion, $\pi:\mathcal{D}\mapsto X$ is the projection of $\mathcal{D}$ on $X$ and $\pi_{k}^{j}:\mathcal{D}_{k}^{(j)}\mapsto X_{k}^{(j)}$ is the projection of $\mathcal{D}_{k}^{(j)}$ on $X_{k}^{(j)}$, there exists a unique graph morphism $w:X_{k}^{(j)}\mapsto X$ such that
$\pi\circ i_{k}^{(j)}=w\circ\pi_{k}^{(j)}.$
@@
Without loss of generality, we can assume that the $r$-strip $S$ is delimitated by straight lines. In particular we can make a reparametrization to ensure that the curve $b$ corresponds to a straight line in the $a$-direction.. For all $j$, consider the strips $\mathcal{B}_{n+j}$ as defined in (4.1) and take $S_{j}=\mathcal{B}_{n+j}$. Observe that for $n$ large enough $S_{j}\subset U\subset S$ and by construction $S_{j}\cap S_{j}^{\prime}=\emptyset$ for all $j\neq j^{\prime}$. Apply Section $4$ and Section $5$ to ensure that $\mathcal{B}_{n+j}$ is indeed a $n+j$- strip.
The descendants are the pairs $\left\{b_{k}^{(j)},U_{k}^{(j)}\right\}$ with $k\in\left\{1,\dots,n\right\}$ where $b_{k}^{(j)}$ is the curve of secondary tangencies as constructed in Proposition 4 and $U_{k}^{(j)}$ is a neighborhood of $b_{k}^{(j)}$ as built in Lemma 17. In particular the family $F_{|U_{k}^{(j)}}$ is an unfolding of the homoclinic tangency $b_{k}^{(j)}$.
Given a descendent of $S_{j}$, $\left\{b_{k}^{(j)},U_{k}^{(j)}\right\}$. Observe that $a_{n+j},b_{k}^{(j)}\in S_{j}=\mathcal{B}_{n+j}$ and the diameter of $\mathcal{B}_{n+j}$ is smaller than the distance between $a_{n}$ and the previous tangency $b$ that is indeed of then order of $1/\mu^{n+j}$, see 2.14. This proves $(iii)$. Condition $(iv)$ on the derivative comes from 5.4 and 5.6. Condition $(v)$ and the following statement are consequence of the construction. ∎
Remark 14.
@@
Observe that, using the notation as in the previous proposition and by using the construction in Subsection 4.1, there exist $n_{k}^{(j)},\underline{\ell}_{k}^{(j)},\overline{\ell}_{k}^{(j)},\theta_{k}^{(j)},n_{0,k}^{(j)}$ such that
$\underline{w}\left(\omega_{k}^{(j)}\right)=\left(q,{q}^{\underline{\ell}_{k}^{(j)}}{c}{q}^{\theta_{k}^{(j)}n_{k}^{(j)}}{c}{q}^{n_{k}^{(j)}}{c}{q}^{n_{0,k}^{(j)}}{h}{q}^{\overline{\ell}_{k}^{(j)}},{q}^{\underline{\ell}_{k}^{(j)}}{h}{q}^{\overline{\ell}_{k}^{(j)}}\right).$

### v3

@@ 6.2 Newhouse phenomenon: infinitely many sinks
For the last property, let $(t^{*},0)$ such that ${\log{\lambda(t^{*},0)^{-1}}}/{\log\mu(t^{*},0)}$ is the maximum of ${\log{\lambda(t,0)^{-1}}}/{\log\mu(t,0)}$. Consider a sequence of first generation Newhouse boxes $\mathcal{P}_{n,n_{0}}\in\mathfrak{N_{1}}$ accumulating at $(t^{*},0)$. This is possible because of Lemma 6.2.5. Choose $\epsilon>0$ and let $n$ be maximal such that $\epsilon\leq{\epsilon_{0}}/{\mu(t^{*},a^{*})^{2n}}$. Because (6.2.2), (6.2.6) and the fact that the vertical size of $\mathcal{P}_{n,n_{0}}$ is ${\epsilon_{0}}/{\mu(t^{*},a^{*})^{2n}}$, we need at least ${K}/{\lambda(t^{*},a^{*})^{\theta n}}$ balls of radius $\epsilon$ to cover $\overline{NH}\cap\mathcal{P}_{n,n_{0}}$. As consequence
@@
and $MD(NH)\geq{2}/{3}$, where we used (2.3.2). ∎
Remark 6.2.8.
@@ 7 Dynamical Partitions, Directed Graphs and Strips
In this section, we introduce the objects and the induction step (see Proposition 7.4.1) that will allow to build later the main player of this work: the lamination and the invariant Cantor set.
@@ 7.1 Dynamical Partitions
We give now the definition of a dynamical partition associated to a map. The reader can refer to Figure 9. The definition models the combinatorics as described in Subsection 4.1.
@@
Recall the rescaling done in Subsection 2.1.1 while introducing our unfoldings and let $Q$ be the domain of linearization. We denote by $W^{u}_{\text{loc}}(p)$ the connected component of $W^{u}(p)\cap Q$ and by $W^{s}_{\text{loc}}(p)$ the connected component of $W^{s}(p)\cap Q$ both containing $p$.
Definition 7.1.1.
@@
    • $C^{N}\subset Q\setminus f(Q)$,
    • $C^{0}\cap f\left(W^{u}_{\text{loc}}(p)\right)\neq\emptyset$,
    • $C^{N}\cap W^{s}_{\text{loc}}(p)\neq\emptyset$.
@@
    • $H^{M}\subset Q\setminus f(Q)$,
    • $H^{0}\cap f\left(W^{u}_{\text{loc}}(p)\right)\neq\emptyset$,
    • $H^{M}\cap W^{s}_{\text{loc}}(p)\neq\emptyset$,
    • $f^{M}\left(H_{0}\cap f\left(W^{u}_{\text{loc}}(p)\right)\right)$ intersects transversally $W^{u}_{\text{loc}}(p)$ in the point $q_{2}$.
Definition 7.1.2.
@@
where diam(U) is the diameter of $U$. Moreover we denote by $\mathcal{D}$ the union of elements in $\mathfrak{D}$. Namely,
@@ 7.2 Directed Graphs
The shortest path from the vertex $v_{1}$ to the vertex $v_{2}$ is denoted by $\overrightarrow{v_{1}v_{2}}$.
Remark 7.2.2.
Observe that $q$ has three ingoing and outgoing edges. All other vertices have exactly one ingoing and one outgoing edge. In particular each vertex other than $q$ has a unique predecessor and a unique successor.
Let $X$ be the $(1,N,M)$ directed graph. A loop $\lambda$ is a path in $X$ which starts and ends in $q$. The loops $\lambda^{q}$, $\lambda^{h}$ and $\lambda^{c}$ are called the prime loops of $X$, respectively $q$-loop, $h$-loop and $c$-loop. Observe that a loop $\lambda$ is a concatenation of prime loops $\gamma_{i}$, namely,
@@
Given $N,M,\tilde{N},\tilde{M}$ positive natural numbers, consider $the (1,N~,M~)(1,\tilde{N},\tilde{M}) directed graph X~\tilde{X}$ and assume there is a graph morphism
@@
In the following lemma we introduce the projection of $D$ onto $X$.
Lemma 7.2.3.
@@ 7.3 Strips
We introduce strips in the plane that will enable us later to construct, inductively, the object of primary interest to us: the lamination. Observe that the strip are an abstract generalization of the $\mathcal{B}_{n}$ strips built earlier to contain the curves of secondary tangencies. We start by defining $r$-tangency strips, with $r$ a positive number; see Figure 11.
Definition 7.3.1.
@@
    $\left|s_{l}(t)-s_{u}(t)\right|\leq C\rho^{r},$
    where $0<\rho<1$, $\rho<\max_{t}\left\{\frac{1}{\mu(t)}\right\}$, and $C$ is a positive constant.
  • $S$ contains $r$ curves $\left\{b_{k}\right\}_{k=1}^{r}$, which are graph of functions $b_{k}:[-t_{0},t_{0}]\mapsto\mathbb{R}$ for all $k$, such that the following holds:
@@
Remark 7.3.2.
Let $f\in S$, then $f$ has a $(1,N,M)$ dynamical partition. We identify $f$ with its restriction to this dynamical partition. If in particular $f\in U_{k}$ for some $k$, then it has a finer dynamical partition $(1,N_{k},M_{k})$ which describes the dynamics on a smaller scale. For this reason we call the restriction of $f$ to the finer dynamical partition the renormalization of $f$.
@@ 7.4 The induction step
Proposition 7.4.1.
Let $S$ be a $r$-tangency strip and let $\left\{b,U\right\}$ be a $(1,N,M)$ descendant of $S$. Then there exist $C>0$ and $n\in\mathbb{N}$, $n>>r$ large enough and a countable sequence of regions $\left\{S_{j}\right\}_{j=1}^{\infty}\subset U\subset S$ satisfying the following:
@@
    $\forall j$, $S_{j}$ is a $(\sqrt{n+j})$-tangency strip,
@@
Moreover, for all $j$ and for all $k\in\left\{1,\dots,\sqrt{n+j}\right\}$ let $\left\{b_{k}^{(j)},U_{k}^{(j)}\right\}$ be a $\left(1,N_{k}^{(j)},M_{k}^{(j)}\right)$ descendant of $S_{j}$. Then the following holds:
@@
    $\forall t\in\left[-t_{0},t_{0}\right]$,
    $\left|b_{k}^{(j)}(t)-b(t)\right|\leq\frac{C}{\mu(t)^{n+j}},$
@@
    $\forall t\in\left[-t_{0},t_{0}\right]$,
    $\left|\frac{db_{k}^{(j)}}{dt}(t)-\frac{db}{dt}(t)\right|\leq C\frac{n+j}{\left(\lambda(t)^{\theta}\mu(t)^{2}\right)^{n+j}},$ (7.4.2)
@@
    $\forall f\in U_{k}^{(j)}\subset U$, let $\mathfrak{D}^{(j)}_{k}(f)$ be the $\left(1,N_{k}^{(j)},M_{k}^{(j)}\right)$ dynamical partition of $f$ and let $\mathfrak{D}(f)$ be the $\left(1,N,M\right)$ dynamical partition of $f$. Then $\mathfrak{D}^{(j)}_{k}(f)$ is a refinement of $\mathfrak{D}(f)$ and
    $\text{mesh}\left(\mathfrak{D}^{(j)}_{k}(f)\right)<\frac{1}{2}\cdot\text{mesh}\left(\mathfrak{D}(f)\right).$
Furthermore, if $X$ is the $\left(1,N,M\right)$ directed graph, $X_{k}^{(j)}$ is the $\left(1,N_{k}^{(j)},M_{k}^{(j)}\right)$ directed graph, $i:\mathcal{D}_{k}^{(j)}\mapsto\mathcal{D}$ is the inclusion, $\pi:\mathcal{D}\mapsto X$ is the projection of $\mathcal{D}$ on $X$ and $\pi_{k}^{j}:\mathcal{D}_{k}^{(j)}\mapsto X_{k}^{(j)}$ is the projection of $\mathcal{D}_{k}^{(j)}$ on $X_{k}^{(j)}$, there exists a unique graph morphism $w_{k}^{(j)}:X_{k}^{(j)}\mapsto X$ such that
$\pi\circ i_{k}^{(j)}=w_{k}^{(j)}\circ\pi_{k}^{(j)}.$
@@
Without loss of generality, we can assume that the $r$-strip $S$ is bounded by straight lines. In particular, we can make a reparametrization to ensure that the curve $b$ corresponds to a straight line in the $t$-direction. For all $j$, consider the strips $\mathcal{B}_{n+j}$ as defined in (4.0.2) and take $S_{j}=\mathcal{B}_{n+j}$. Observe that for $n$ large enough $S_{j}\subset U\subset S$ and by construction $S_{j}\cap S_{j^{\prime}}=\emptyset$ for all $j\neq j^{\prime}$. Moreover Section $4$ and Section $5$ ensure also that $\mathcal{B}_{n+j}$ contains $\sqrt{n+j}$ secondary tangency curves. The descendants are the pairs $\left\{b_{k}^{(j)},U_{k}^{(j)}\right\}$ with $k\in\left\{1,\dots,\sqrt{n+j}\right\}$ where $b_{k}^{(j)}$ is the curve of secondary tangencies as constructed in Proposition 5.3.1 and $U_{k}^{(j)}$ is a neighborhood of $b_{k}^{(j)}$ as built in Lemma 5.3.3. In particular, the family $F_{|U_{k}^{(j)}}$ is an unfolding of the homoclinic tangency $b_{k}^{(j)}$. Given a descendent of $S_{j}$, $\left\{b_{k}^{(j)},U_{k}^{(j)}\right\}$, observe that $a_{n+j},b_{k}^{(j)}\subset S_{j}=\mathcal{B}_{n+j}$ and the width of $\mathcal{B}_{n+j}$ is smaller than the distance between $a_{n}$ and the previous tangency $b$ that is indeed of then order of $1/\mu^{n+j}$, see (2.2.10). This proves $(iii)$. Condition $(iv)$ on the derivative comes from (5.1.7). The final step is to construct the corresponding dynamical partitions. Let $f\in b_{k}^{(j)}$. In particular there is a point $m_{1}\in W^{u}_{\text{loc}}(p)$ and a time $s$ such that at $f^{s+M}(m_{1})$ there is a tangency of $W^{s}(p)\cap W^{u}_{\text{loc}}(p)$, see Section $4$. Let $Q_{k}^{(j)}\subset Q$ such that
$\overline{Q_{k}^{(j)}}\cap\left\{m_{1},f(m_{1}),\dots,f^{s+M}(m_{1})\right\}=\emptyset.$
Let $\overline{\ell}_{k}^{(j)}$ be minimal such that
$f^{-\overline{\ell}_{k}^{(j)}+1}(q_{2}),f^{-\overline{\ell}_{k}^{(j)}+1}(m_{1})\in\overline{Q_{k}^{(j)}},$
and let $\underline{\ell}_{k}^{(j)}$ be minimal such that
$f^{\underline{\ell}_{k}^{(j)}}(f^{M}(q_{2})),f^{\underline{\ell}_{k}^{(j)}}(f^{s+M}(m_{1})).$
Let $(C^{0})^{(j)}_{k}$ be a small enough neighborhood of $f^{-\overline{\ell}_{k}^{(j)}+1}(m_{1})$ and let $(H^{0})^{(j)}_{k}$ be a small enough neighborhood of $f^{-\overline{\ell}_{k}^{(j)}+1}(q_{2})$ such that
$\mathfrak{D}_{k}^{(j)}(f)$ $=$ $\left\{\overline{Q}_{k}^{(j)}\right\}\cup$
$\left\{\overline{f^{i}\left((C^{0})^{(j)}_{k}\right)}\left|\right.i\in\left\{0,\dots,\overline{\ell}_{k}^{(j)}+s+M+\underline{\ell}_{k}^{(j)}-1\right\}\right\}\cup$
$\left\{\overline{f^{i}\left((H^{0})^{(j)}_{k}\right)}\left|\right.i\in\left\{0,\dots,\overline{\ell}_{k}^{(j)}+M+\underline{\ell}_{k}^{(j)}-1\right\}\right\}$
forms a pairwise disjoint collection,
$\text{mesh}\left(\mathfrak{D}^{(j)}_{k}(f)\right)<\frac{1}{3}\cdot\text{mesh}\left(\mathfrak{D}(f)\right).$
and
$\text{dist}\left({f^{\overline{\ell}_{k}^{(j)}+s}\left((C^{0})^{(j)}_{k}\right)},q_{2}\right)<\lambda^{n_{0,k}^{(j)}/2},$ (7.4.3)
where $n_{0,k}^{(j)}$ is defined in Subsection 4.1. Observe that the construction of $\mathfrak{D}^{(j)}_{k}(f)$ can be made continuous along the curve $b_{k}^{(j)}$. Moreover by restricting the neighborhood $U_{k}^{(j)}$ we can also make the construction of $\mathfrak{D}^{(j)}_{k}(f)$ varying continuously over the neighborhood still satisfying condition (7.4.3). In particular, $\mathfrak{D}^{(j)}_{k}(f)$ is a dynamical partition with $N^{(j)}_{k}=\overline{\ell}_{k}^{(j)}+s+\underline{\ell}_{k}^{(j)}-1$ and $M^{(j)}_{k}=\overline{\ell}_{k}^{(j)}+M+\underline{\ell}_{k}^{(j)}-1$. This proves condition $(v)$ and the last statement is consequence of this construction. ∎
Remark 7.4.4.
@@
Observe that, using the notation as in the previous proposition and by using the construction in Subsection 4.1, there exist $n_{k}^{(j)},\underline{\ell}_{k}^{(j)},\overline{\ell}_{k}^{(j)},n_{0,k}^{(j)}$ such that
$\underline{\omega}\left(w_{k}^{(j)}\right)=\left(q,{q}^{\underline{\ell}_{k}^{(j)}}{h}{q}^{n_{0,k}^{(j)}}{c}{q}^{n_{k}^{(j)}}{c}{q}^{\theta n_{k}^{(j)}}{c}{q}^{\overline{\ell}_{k}^{(j)}},{q}^{\underline{\ell}_{k}^{(j)}}{h}{q}^{\overline{\ell}_{k}^{(j)}}\right).$